Find the sums of the following fractions:
1. 1/4+1/8.
2. 1/2+3/10.
3. 5/12+1/4.
4. 2/5+4/15.
5. 2/9+2/3.
6. 2/3+7/15.
7. 8/21+3/7.
1
Expert's answer
2020-09-21T11:22:42-0400

1. 1/4+1/8 = (2+1)/8

= 3/8

2. 1/2+3/10 = (5+3)/10

= 8/10

= 4/5

3. 5/12+1/4 = (5+3)/12

= 8/12

= 2/3

4. 2/5+4/15 = (6+4)/15

= 10/15

= 2/3

5. 2/9+2/3 = (2+6)/9

= 8/9

6. 2/3+7/15 = (10+7)/15

= 17/15

7. 8/21+3/7 = (8+9)/21

= 17/21